Test II – Computation of Demand, Supply, Income & Cross Elasticity: Show your solution and box your final answers. Round off to two decimal points.

1) Leeroy Jenkins is engaged in the sale of high-classed furniture. He noticed that one particular table that costs P1000 sold 500 pcs last month. Due to oil price hike of 5%, all the prices of other commodities in market also increased by 10%. Since Leeroy is a cautious businessman, he also increased the price of his table to P1500. As such, he noticed that only 200 pcs of his table were sold this month. Much to his dismay, he could only think of one thing, he wants to try out for an NBA team. Compute for demand elasticity of the product. Is it elastic, inelastic or unitary? Write it beside your demand elasticity coefficient.
1
Expert's answer
2016-07-24T11:47:02-0400
If P1 = 1000 and Q1 = 500 pcs, but P2 = 1500 and Q2 = 200 pcs, then the demand elasticity of the product is:
Ed = (200 - 500)/(1500 - 1000)*(1500 + 1000)/(200 + 500) = -300/500*2500/700 = -15/7 = -2.14, so the demand is elastic.